### v1

# Quantum Otto engine mimicking Carnot near pseudotransitions in the 1D extended Hubbard model in the atomic limit

Abstract
The one-dimensional extended Hubbard model (EHM) in the atomic limit has recently been found to exhibit a curious thermal pseudo-transition behavior, which closely resembles first and second-order thermal phase transitions. This phenomenon, occurring at half-filling, is influenced by the quantum phase transition between the alternating pair (AP) and paramagnetic (PM) phases at zero temperature. In this study, we leverage this anomalous behavior to investigate the performance of quantum many-body machines, using the EHM as the working substance. Our analysis reveals that the quantum Otto engine, when operating in the anomalous region, closely mimics the ideal Carnot engine. In this region, both the work output and thermal efficiency of the Otto engine increase, approaching the performance of a Carnot engine. This highlights the potential of many-body systems, such as the EHM, in enhancing quantum thermodynamic performance. Our findings demonstrate that, although the second law of thermodynamics prevents engines from surpassing Carnot efficiency, the Otto engine can operate remarkably close to this limit in the anomalous region, offering insights into new directions for future research on quantum thermodynamic cycles and working substances.
I Introduction
In the past two decades, significant efforts have been made to explore thermodynamic processes where quantum features of matter are crucial. This research has successfully combined theoretical proposals [1] with experimental evidence [2], bridging the gap between quantum thermodynamics concepts and practical applications. Quantum heat engines, like Carnot and Otto, have been studied, with their characteristics compared to classical counterparts. Additionally, a quantum heat engine model has been analyzed for efficiency and irreversibility [3, 4]. Quantum thermodynamics offers a framework for understanding and optimizing engines and refrigerators [5]. The study of quantum heat engines based on harmonic oscillators has enhanced our understanding of optimization and irreversibility [6, 7].
Recent investigations have advanced quantum Carnot, Otto, and Stirling engines, surpassing previous limitations in achieving Carnot efficiency with quantum and nanoscale heat engines. These include achieving Carnot efficiency through semi-local thermal operations with dual bath interactions [8]. Classical magnetic Otto cycles outperformed quantum ones in work extraction due to thermodynamic equilibrium [9]. Energy quantization effects in the Otto cycle were found to enhance performance, with experimental proposals using trapped ions [10]. Several works have explored theoretical dimer systems as working substances, including quantum engines, refrigerators, and heaters within a quantum Otto cycle with coupled spins [11, 12, 13], and including Dzyaloshinskii–Moriya interactions [14]. Though single-spin and spin lattice systems have been extensively studied [15, 16], trimer systems have been rarely explored, with one notable quantum Otto cycle analysis [17].
Several quantum machines have been explored using one-dimensional models. In this context, a quantum Otto heat engine based on a multiferroic chain as the working substance was investigated in [18]. Halpern et al. [19] proposed harnessing many-body localization for thermodynamic tasks, designing an Otto engine with mesoscale parallel reliability. In [20], the thermodynamic cycle of an Otto engine with a quantum Ising chain as the working substance was examined. This cycle involves sweeps of the transverse magnetic field in thermal isolation, alternated with thermalization strokes involving reservoirs at different temperatures. The performance of quantum thermal machines using many-body quantum systems or spin-chain models, periodically connected to external baths via local couplings, was discussed in [21]. More recently, the spin-chain-star shape model was analyzed as the working substance of a quantum Otto cycle in [22]. Additionally, researchers have studied a finite-time quantum Otto cycle in a 1D Bose gas, analyzing efficiency-power trade-offs [23]. Nautiyal et al. [24] examined an imperfect quantum Otto engine, evaluating work, power, and optimal operating points under less ideal conditions.
The Hubbard model [25] is fundamental in studying strongly correlated electrons, particularly where interactions are crucial. The one-dimensional extended Hubbard model (EHM), which includes nearest-neighbor interaction energy, is widely studied for its relevance to materials like carbon nanotubes and organic conductors[26]. It serves as a prototype for strongly correlated electron systems [27]. Tsuchiizu and Furusaki [28] revisited its ground-state phase diagram, revealing new phases and transitions. Glocke et al. [29] studied thermodynamics, highlighting phase transitions through compressibility and susceptibility. The phase diagram of the one-dimensional $U-V$ model at quarter-filling shows a transition from Luttinger liquid to charge density wave insulator, influenced by superconducting or spin density wave fluctuations[30]. In the 1970s, researchers [31, 32, 33] used the transfer matrix to study EHM thermodynamics properties such as specific heat, magnetic susceptibility, and density correlations at half and quarter fillings with infinite repulsion. Later, Mancini and Mancini[34, 35] identified four phases, charge orderings, and critical polarization fields [36].
Quasi-one-dimensional (quasi-1D) materials exhibit fascinating electronic properties due to reduced dimensionality. For example, $\beta\text{\textquoteright}-\mathrm{EtMe}{{}_{3}}\mathrm{Sb[Pd(dmit}){{}_{2}}]{{}_{2}}$ is a molecular triangular lattice system proposed as a quantum spin liquid candidate with quasi-1D spin dynamics [37]. Organic charge-transfer salts like $\mathrm{(TMTTF)}{{}_{2}}\mathrm{X}$ and $\mathrm{(TMTSF)}{{}_{2}}\mathrm{X}$ show strong electron correlations and charge-ordering phenomena[38]. The TTF finite-energy spectral features in photoemission of TTF–TCNQ are explained by a Hubbard-chain description, revealing that $U$ is larger for TTF than TCNQ[39]. One-dimensional electronic transport at the organic charge-transfer interfaces under high pressure[40]. Layered systems like $\mathrm{SrCuO}{{}_{2}}$ reveal spinon-holon separation, further emphasizing low-dimensional physics[41]. These materials collectively highlight dimensionality’s role in emergent quantum phenomena. Additionally, angle-resolved photoemission spectroscopy (ARPES) studies on 1D extended Hubbard model systems have provided valuable insights into electron-phonon coupling effects, further validating theoretical predictions[42].
The present work is organized as follows: In Sec. 2, we summarize the thermodynamics of the one-dimensional EHM in the atomic limit, particularly around anomalous region. In Sec. 3, we discuss thermodynamic process of quantum many body machines. In Sec. 4, we discuss our results regarding quantum machine operation modes influenced by pseudo-transitions. Finally, in Sec. 5, we present our conclusions and perspectives.
II 1D EHM in the atomic Limit and its thermodynamics
Therefore, we summarize the most relevant results obtained in reference [43]. Despite the simplest version of the EHM has been extensively studied across various physical systems, our focus is on a one-dimensional EHM in the atomic limit, where the hopping term is neglected, resulting in the following Hamiltonian
$\boldsymbol{H}=$ $\sum_{i=1}^{N}\left[U\boldsymbol{n}_{i,\uparrow}\boldsymbol{n}_{i,\downarrow}+V\boldsymbol{n}_{i}\boldsymbol{n}_{i+1}\right.$
$\left.-\mu(\boldsymbol{n}_{i,\uparrow}+\boldsymbol{n}_{i,\downarrow})-h(\boldsymbol{n}_{i,\uparrow}-\boldsymbol{n}_{i,\downarrow})\right].$ (1)
Here, $U$ denotes the on-site Coulomb interaction between electrons occupying the same lattice site, $V$ represents the interaction between electrons on neighboring sites, $\mu$ is the chemical potential that governs the electron filling of the lattice, and $\boldsymbol{n}_{i,\sigma}$ represents the number operator at site $i$ with spin $\sigma=\{\uparrow,\downarrow\}$ , where $\boldsymbol{n}_{i}=\boldsymbol{n}_{i,\uparrow}+\boldsymbol{n}_{i,\downarrow}$. The last term accounts for the contribution of the external magnetic field $h$.
II.1 Transfer matrix results
[…]
The model can be solved using the transfer matrix technique[32, 43], with the transfer matrix given by
$\mathbf{W}=\left[\begin{array}[]{cccc}1&yw_{0,1}&y^{-1}w_{0,1}&w_{0,2}\\ yw_{0,1}&y^{2}w_{1,1}&w_{1,1}&yw_{1,2}\\ y^{-1}w_{0,1}&w_{1,1}&y^{-2}w_{1,1}&y^{-1}w_{1,2}\\ w_{0,2}&yw_{1,2}&y^{-1}w_{1,2}&w_{2,2}\end{array}\right],$ (2)
where $w_{0,1}={\rm e}^{\beta\mu/2}$, $w_{0,2}={\rm e}^{\beta(\mu-U/2)}$, $w_{1,1}={\rm e}^{\beta(\mu-V)}$, $w_{1,2}={\rm e}^{\beta(3\mu/2-2V-U/2)}$, $w_{2,2}={\rm e}^{\beta(2\mu-4V-U)}$, $y={\rm e}^{\beta h/2}$, and $\beta=1/(k_{B}T)$.
The eigenvalues of the transfer matrix consist of one trivial null eigenvalue, while the remaining eigenvalues are determined by the following cubic secular equation
$\lambda^{3}+a_{2}\lambda^{2}+a_{1}\lambda+a_{0}=0,$ (3)
where the coefficients are given by:
$a_{0}=$ $z\left(w_{1,2}^{2}+w_{0,2}^{2}w_{1,1}+w_{0,1}^{2}w_{2,2}\right)$
$-z\left(w_{1,1}w_{2,2}-2w_{0,2}w_{0,1}w_{1,2}\right),$
$a_{1}=$ $z\left(w_{1,1}+w_{1,1}w_{2,2}-w_{0,1}^{2}-w_{1,2}^{2}\right)+w_{2,2}-w_{0,2}^{2},$
$a_{2}=$ $-1-zw_{1,1}-w_{2,2}.$ (4)
Therefore, the roots of the algebraic cubic equation (3) can be conveniently expressed using trigonometric functions as follows
$\lambda_{j}=2\sqrt{Q}\cos\left(\tfrac{\phi-2\pi j}{3}\right)-\frac{1}{3}a_{2},\quad j=0,1,2,$ (5)
where
$\phi$ $=\arccos\left(\tfrac{R}{\sqrt{Q^{3}}}\right),$ (6)
$Q$ $={\displaystyle\frac{a_{2}^{2}-3a_{1}}{9},}$ (7)
$R$ $={\displaystyle\frac{9a_{1}a_{2}-27a_{0}-2a_{2}^{3}}{54}.}$ (8)
Additional details regarding the solution, with the largest eigenvalue $\lambda_{0}$, can be found in reference [43].
II.2 Thermodynamic quantities
To analyze thermodynamic quantities, we use the grand partition function for a chain of $N$ sites:
$\Xi(T,\mu,h,N)=\lambda_{0}^{N}+\lambda_{1}^{N}+\lambda_{2}^{N},$ (9)
where the eigenvalues are ordered as $\lambda_{0}>\lambda_{1}>\lambda_{2}$. The grand potential per site in the thermodynamic limit $N\to\infty$, is determined by the largest eigenvalue of the transfer matrix:
$\Omega(T,\mu,h)=$ $-k_{B}T\ln\lambda_{0}.$ (10)
We define some thermodynamic quantities such as entropy ${\cal S}=-\frac{\partial\Omega}{\partial T}$, enthalpy $\mathcal{E}=k_{B}T^{2}\frac{\partial\ln\left(\lambda_{0}\right)}{\partial T}$, magnetization $M=-\frac{\partial\Omega}{\partial h}$, and electron density $\rho=-\frac{\partial\Omega}{\partial\mu}$.
II.3 Frustrated and alternating pair phase transition
Here we focus in two ground state phases at half-filling band $\rho=1$. The first phase is the frustrated phase $FR_{2}$ given by
$|FR_{2}\rangle=$ $_{j=1}^{N}|\updownarrow\rangle,$
[…]
Its ground-state energy is
$E_{FR_{2}}=$ $V-h-\mu.$ (12)
The corresponding residual entropy is $\mathcal{S}=k_{B}\ln(2)$. This state becomes non-frustrated when $h\neq 0$, aligning with the magnetic field.
The other phase is the alternating pair phase ($AP$) expressed by
[…]
Its ground-state energy is
$E_{AP}=$ $\frac{U}{2}-\mu.$ (14)
The AP phase has no residual entropy and can be identified also as a charge density wave (CDW)[29].
The ground-state phase boundary between both phases is given by $h=V-\frac{U}{2}$.
II.4 Pseudo-critical temperature condition
Recent studies on one-dimensional models with short-range interactions have revealed pseudo-transitions [44], which can be analyzed through spin correlations [45]. These pseudo-transitions have been observed in several models, including the Ising diamond[46, 47], ladder [48, 49, 50], Ising-Heisenberg diamond [51, 52], double-tetrahedral[53], and Ising-Heisenberg ladder models [54].
In the context of ground-state phase transitions, we summarize the thermal behavior in the low-temperature region, as explored in Ref. [43].
At low temperatures, an anomalous behavior is observed at a characteristic temperature, referred to as the pseudo-critical temperature $T_{p}$ [43], which satisfies the following relation
$w_{1,1}z-w_{0,2}=0.$ (15)
Explicitly, with $w_{0,2}={\rm e}^{\beta(\mu-U/2)}$, $w_{1,1}={\rm e}^{\beta(\mu-V)}$, and $z=2\cosh(\beta h)$, we have the following expression
$2\cosh(\beta_{p}h)={\rm e}^{\beta_{p}(V-U/2)}.$ (16)
Particularly, for $0<h\lesssim V-\frac{U}{2}$, an anomalous behavior arises, indicating a pseudo-transition at temperature $T_{p}$.
Although Eq.(16) is independent of the chemical potential $\mu$, this condition is valid only within a limited interval, $21<\mu<$ 2. For details, refer to the phase diagram in Ref. [43].
The sharp, yet continuous, entropy jump near the pseudo-critical temperature $T_{p}$, although not signaling a true phase transition, reveals a rapid microscopic reconfiguration in the system. This enhanced response suggests that even minor temperature variations can lead to substantial changes in the working substance’s thermal properties. Notably, $T_{p}$ is unrelated to a true critical temperature, yet studies [48, 55, 56] suggest universality in pseudo-critical exponents, adhering to the Rushbrooke inequality. Similar behavior appears in the extended Hubbard, Potts [57], Zimm-Bragg-Potts [58], and spin-pseudospin models [59, 60].
II.5 Magnetic Properties
Here, we examine the magnetic influence near the pseudo-transition. At low temperatures, quantum phase transitions often enhance the magnetocaloric effect, a well-studied phenomenon. Following this perspective, we analyze low-temperature isentropic curves in regions near quantum phase transitions occur at zero temperature.
Figure 1: Isentropic curve around anomalous region, in the plane $h-T$, assuming $U=10$, $V=5.4$, $\mu=15.8$
In Fig. 1, the isentropic curve is illustrated in the $h-T$ plane, considering the following fixed parameters: $U=10$, $V=5.4$, $\mu=15.8$. The isentropic curve exhibits unusual behavior over a certain range of entropy, as shown in the legend. Specifically, for $0.40.3\apprle h<$ 0.4, all isentropic curves are nearly overlapping. This unusual behavior significantly affects the thermal properties and, for lower magnetic fields, the curves remain densely packed, highlighting the influence of the anomalous region. The blue curves delineate the quasi-alternating pair ($qAP$) region for $h\lesssim 0.4$, whereas for $h\gtrsim 0.4$, the area below the blue curve corresponds to the quasi-paramagnetic ($qPM$) region. Additionally, for a null magnetic field and at temperatures roughly around $T\sim 0.5-0.6$, there is evidence of a strong change in entropy, indicating a transition from the $qAP$ to the $qFR_{2}$ region. In Fig. 1, below blue curve ($T\lesssim 0.01$) basically the entropy is null, see Fig. 2a from another perspective.
Fig. 1 shows unusual thermodynamic behavior, with nearly overlapping isentropic curves indicating an enhanced magnetocaloric effect in the pseudo-critical region. Despite the absence of singularities, sharp entropy redistribution occurs, potentially benefiting quantum thermodynamic cycles by enabling efficient thermal control without fine-tuned critical conditions.
For this purpose, in fig.2a, the entropy is shown as a function of temperature, assuming $U=10$, $V=5.3$ and $\mu=15.8.$ For a zero magnetic field and $h_{1}=0.29$, we observe a typical abrupt yet continuous change in entropy. This jump occurs for $h_{0}=0$ at $T\sim 0.2$ and for $h_{1}=0.29$, the jump occurs at $T\sim 0.43$. This continuous jump mimics a first order phase transition; however, it is important to note that no true phase transition takes place.
In panel (b), we illustrate the variation of entropy $\Delta\mathcal{S}=\mathcal{S}(T,h_{1})-\mathcal{S}(T,0)$ as a function of temperature. We assume $U=10$, $\mu=15.8$ and $V=5.3$, with $h_{1}=0.29$ and $h_{0}=0$. The entropy variation $\Delta\mathcal{S}$ is represented by the violet curve. Below the jump, $\Delta\mathcal{S}$ is very small, and a strong jump occurs at $T\sim 0.2$, leading to a large $\Delta\mathcal{S}$, followed by a strong change that results in a negative value at $T\sim 0.43$, after which it increases gradually. A similar behavior is observed for $U=10$, $\mu=15.8$, and $V=5.4$, with $h_{1}=0.38$ and $h_{0}=0$, where $\Delta\mathcal{S}$ is represented by the green curve. The jump in $\Delta\mathcal{S}$ is again small, with a strong jump occurring at $T\sim 0.29$, leading to a large $\Delta\mathcal{S}$, followed by a strong change that results in a negative value at $T\sim 0.56$, and then increasing gradually for higher temperatures.
In panel (c), the adiabatic temperature change for the same set of parameters as in panel (b) is shown, which is given by
$\Delta T_{ad}=-\int_{h_{0}}^{h_{1}}\left(\frac{\partial T}{\partial\mathcal{S}}\right)_{h}\left(\frac{\partial M}{\partial T}\right)_{h}{\rm d}h,$
as a function of temperature. In practice, we can determine $\Delta T_{ad}$ by solving the following equation $\mathcal{S}(T+\Delta T_{ad},h_{1})=\mathcal{S}(T,h_{0})$. To compare our results, we use the same curve colors as in panel (b). The violet curve exhibits a sharp peak at $T\sim 0.2$, which decreases almost linearly, reaching zero around $T\sim 0.43$, followed by a minimum at approximately $T\sim 0.85$, after which increases monotonically. A similar behavior is observed for the green curve, with a sharp peak at $T\sim 0.29$, a minimum at $T\sim 0.95$, and a monotonic increase for higher temperatures.
Panel (d) presents a further analysis of the quantity $\Delta\mathcal{S}\Delta T_{ad}$, which could be particularly relevant for the operation of a Carnot engine. In this analysis, we continue using the same set of parameters established in panel (b). Notably, for temperatures below $T\approx 0.2$ (violet) and $T\approx 0.29$ (green), $\Delta\mathcal{S}\Delta T_{ad}\sim 0$. Then, a significant amount of $\Delta\mathcal{S}\Delta T_{ad}$ appears, vanishing at $T\sim 0.44$ and $T\sim 0.58$, respectively. For higher temperatures, the quantity increases, reaching a maximum at $T\sim 0.84$ and $T\sim 0.91$, respectively. After these peaks, the quantity begins to gradually decrease.
Figure 2: (a) Entropy as a function of temperature $T$, with parameters $U=10$, $V=5.3$ and $\mu=15.8$. (b) Entropy variation $\Delta\mathcal{S}=\mathcal{S}(T,h_{1})-\mathcal{S}(T,0)$ as a function of temperature $T$, with fixed parameters $U=10$, $\mu=15.8$, and $V=5.3$, for $h_{1}=0.29$ (violet curve). Similarly, for $h_{2}=0.38$ (green curve) and fixed parameters $U=10$, $\mu=15.8$, and $V=5.4$. (c) Adiabatic temperature change $\Delta T_{ad}$ as a function of temperature, for the same set of parameters as in (b). (d) The product $\Delta\mathcal{S}\Delta T_{ad}$ as a function of temperature, under the same conditions as in (b).
The entropy variation $\Delta\mathcal{S}$ in Fig. 2b reveals a sharp yet continuous crossover, mimicking a first-order transition without true discontinuities. This pseudo-transition reflects a rapid microscopic reconfiguration without singularities in response functions, ensuring stable behavior across all temperatures. Such controlled entropy changes could be useful for quantum thermodynamic processes.
III Reversible cycle Process
In the following, we proceed to analyze the thermal machine cycles. Particularly, we will focus on the reversible Carnot and Otto engines.
III.1 Work and heat transferred
The grand potential, denoted as $\Omega$ and defined by (10), is a key quantity for analyzing systems in the grand canonical ensemble. This ensemble describes systems where the particle number can fluctuate, and external magnetic field act as control parameters. The first law of thermodynamics for the present system can be expressed as:
${\rm d}{\cal E}=T{\rm d}{\cal S}-M{\rm d}h-\rho{\rm d}\mu,$ (17)
[…]
Here, the quantities ${\cal E}$ (enthalpy), ${\cal S}$ (entropy), $M$ (magnetization), and $\rho$ (particle density) are defined per lattice site to normalize extensive variables in systems with fixed lattice size.
III.2 Operation modes of machine
Considering the one-dimensional EHM in the atomic limit as the working substance, the work is linked to changes in energy levels as the magnetic field $h$ is altered. By adjusting the parameters of the reversible cycle in a 1D EHM in the atomic limit, the total work of the cycle can be reversed in accordance with the second law of thermodynamics, which, as formulated by Clausius, defines operational regions for thermodynamically reversible cycles based on temperature differences ($T_{h}>T_{l}$). In heat engine mode, the system absorbs heat from a hotter bath at $T_{h}$ ($Q_{in}>0$) and releases it to a cooler bath at $T_{l}$ ($Q_{out}<0$), partially converting the heat into work ($W_{net}>0$), as shown in Table 1. In refrigerator mode, the model induces heat flow from the cooler bath at $T_{l}$ to the hotter bath at $T_{h}$, requiring more work input than work output, with net work performed ($W_{net}<0$). As a thermal accelerator, the system uses net work ($W_{net}<0$) to increase heat flow from the hotter to the cooler bath. In heater mode, the net work ($W_{net}<0$) is applied to generate heat flow in both baths, causing the 1D EHM in the atomic limit to release heat in both baths ($Q_{in}<0$ and $Q_{out}<0$). All these processes are summarized in Table 1. In particular, we apply the Carnot and Otto engine processes as detailed in Appendix A.
Table 1: Characteristics of work and heat as allowed by the second law of thermodynamics. The sign ($+$) indicates work done by the system and heat absorbed, while the sign (--) indicates work done on the system and heat released.
Operation mode $W_{net}$ $Q_{in}$ $Q_{out}$ Thermal efficiency
Heat engine $+$ $+$ -- $\eta=\frac{W_{net}}{Q_{in}}$
Refrigerator -- -- $+$ $COP=\frac{Q_{in}}{W_{net}}$
Heater -- -- -- $COP=\frac{Q_{out}}{W_{net}}$
Accelerator -- $+$ -- $COP=\frac{Q_{out}}{W_{net}}$
Fig. 2c shows the adiabatic temperature change $\Delta T_{ad}$, essential for the Otto cycle. The sharp entropy variation in the pseudo-transition region enhances the thermodynamic response, where small parameter changes significantly impact work and efficiency. This suggests that quantum engines could leverage this region for optimal performance without critical divergences. The smooth yet sharp entropy change enables near-reversible cycles, allowing the Otto cycle to closely mimic Carnot efficiency.
III.3 Thermal efficiency
In the heat engine mode, the efficiency of the quantum engine using the 1D EHM in the atomic limit as the working substance is presented in Table 1, where $\eta=\frac{W_{net}}{Q_{in}}<1$. For the other operational modes, the coefficient of performance (COP) is used, as detailed in Table 1. Since COP values are typically greater than one, we introduce an alternative metric,
$\kappa=\frac{COP}{1+COP},$ (18)
which maps the COP to a restricted range of $10<\kappa<$ 1. This metric implies that as COP approaches 00, $\kappa$ also approaches 00, and as COP tends toward infinity, $\kappa$ approaches $1$. Notably, a COP of $1$ corresponds to $\kappa=0.5$. As a result, the thermal efficiency for all operational modes: refrigerator, heater, and accelerator can be consistently defined within this range, where 00 indicates the worst performance and 1 represents the best, as described in Table 1.
For the four-stroke process, the net work performed is given by $W_{net}=W_{ab}+W_{cd}=Q_{bc}+Q_{da}$. Therefore, the efficiency of the Otto engine is expressed as
$\eta_{o}=$ $\frac{W_{net}}{Q_{bc}}=1+\frac{Q_{da}}{Q_{bc}},$ (19)
since the cyclic process occurs around two pseudo-critical temperatures, we express it algebraically in this region to illustrate the efficiency of the Otto engine, as detailed in Appendix B.
From Appendix B, the efficiency of the Otto engine around the pseudo-transition is approximately given by
$\eta_{o}=\eta_{c}-\frac{T_{p_{1}}}{T_{p_{2}}}\left(\frac{\mathcal{S}_{0}+\mathcal{S}_{1}}{2C_{p_{2}}}-\frac{\mathcal{S}_{p_{2}}}{C_{p_{2}}}\right).$ (20)
As expected, the Otto efficiency is never greater than the Carnot efficiency, $\eta_{c}\geqslant\eta_{o}$, since the last term is a positive quantity. However, this term becomes sufficiently small, leading to $\eta_{o}\rightarrow\eta_{c}$.
This is an intriguing result, as it demonstrates that by simply following a four-stroke reversible Otto engine process, one can achieve a near-maximum efficiency comparable to the Carnot efficiency, without requiring any additional modifications to enhance performance.
IV Results of Carnot and Otto machine
The classical Carnot Cycle, regarded as the most efficient thermodynamic cycle, contrasts with the Otto Cycle, which idealizes spark-ignition engines. Recently, Kroetz[61] reexamined this comparison didactically, emphasizing the impracticality of the Carnot cycle in real engines due to its dependency on the compression ratio.
The quantum Carnot and Otto cycles, composed of a four-stroke process, are explored under general conditions detailed in Appendix A, where express the explicit expressions for heat absorbed, heat released and work performed for both engines.
In Fig. 3, the Otto cycle engine is illustrated around the anomalous region [43]. Here, we assume fixed parameters $\mu=15.8$, $U=10$ and $V=5.4$, satisfying the conditions of the anomalous region. In this region, we notice an intriguing behavior of the Otto cycle that mimics the Carnot cycle: the isomagnetic process behaves almost like an isothermal process (vertical line). By “isomagnetic” process we mean a process occurring at a constant magnetic field. The Otto processes, as described in Appendix A, can be performed, for instance, by fixing the following two adiabatic processes at $\mathcal{S}_{0}=0.05$ and $\mathcal{S}_{1}=0.2$, and two isomagnetic processes at $h_{0}=0$ and $h_{1}=0.38$. In Table 2, the cold temperature ($T_{l}$), hot temperature ($T_{h}$), and two temperatures corresponding to the pseudo-critical points ($T_{p_{1}}$ and $T_{p_{2}}$) are reported. Notably, the specific heat at $T_{p_{1}}$ and $T_{p_{2}}$ is considerably larger compared to the specific heat at the cold and hot temperatures.
During the isomagnetic process, the system resembles an isothermal process (represented by an almost vertical line). Consequently, this cyclic process (green cycle) approaches the behavior of a Carnot engine, becoming nearly rectangular in shape. The algebraic expression for the efficiency approximation of this process is detailed in Appendix B as an illustrative example. However, all figures were generated from the exact solution without any approximations.
Table 2: Illustrative temperatures, entropies and specific heat, for the 1D EHM in the atomic limit, assuming the parameters: $\mu=15.8$ and $U=10$.
Temperature $h$ $V$ Entropy Specific heat
$T_{l}$ $0.2883031789$ $0.38$ $5.4$ $0.05$ $64.5881093600$
$T_{p_{1}}$ $0.2885145480$ $0.38$ $5.4$ $0.1231215757$ $123.9985479937$
$T^{\prime}_{l}$ $0.2887433779$ $0.38$ $5.4$ $0.2$ $59.3619717363$
$T^{\prime}_{h}$ $0.5416370179$ 00 $5.4$ $0.05$ $3.6339831194$
$T_{p_{2}}$ $0.5730760977$ 00 $5.4$ $0.3522945133$ $9.9532683846$
$T_{h}$ $0.5634744399$ 00 $5.4$ $0.2$ $7.3916139289$
$T_{l}$ $0.1964545082$ $0.29$ $5.3$ $0.05$ $2888.0095116037$
$T_{p_{1}}$ $0.1964570150$ $0.29$ $5.3$ $0.0987144858$ $4389.4535640584$
$T^{\prime}_{l}$ $0.1976924508$ $0.29$ $5.3$ $0.2$ $0.4129700921$
$T^{\prime}_{h}$ $0.4260966401$ 00 $5.3$ $0.05$ $5.3306912666$
$T_{p_{2}}$ $0.4325649629$ 00 $5.3$ $0.2722766206$ $38.0622167546$
$T_{h}$ $0.4307158885$ 00 $5.3$ $0.2$ $28.2638604911$
Figure 3: The Otto cycle, which mimics the Carnot cycle, is executed by the 1D EHM in the atomic limit using the parameters listed in Table 2, where $\mathcal{S}_{0}=0.05$ and $\mathcal{S}_{1}=0.2$.
[…]
Figure 4: Heat absorbed ($Q_{in}$), Released heat ($Q_{out}$), and Work performed ($W_{net}$) as a function of Magnetic field, assuming $U=10$, $V=5.4$, $\mu=15.8$. Particularly here we fixed the $T_{h}=0.56$ and $T_{l}=0.28$.
In Fig. 5, we illustrate the thermal efficiency under the same conditions as in Fig. 4 to compare the thermal efficiencies of the Carnot and Otto machines, represented by the red and blue curves, respectively. For $h>0.46$, there is a maximum work output, which occurs roughly at $h=0.70$, as shown in Fig. 4. This point is marked in Fig. 5 by a small line, and the corresponding horizontal line (representing the same efficiency) intersects the first region of the thermal engine at approximately $h\approx 0.36$. Therefore, the left side of the violet region corresponds to a range where the efficiencies of both thermal engines are equivalent. In this violet region, the work performed by the Otto engine and its thermal efficiency increase, approaching the work and efficiency of the Carnot engine.
This is an interesting result since our working substance (1D EHM in the atomic limit) becomes highly efficient in the Otto engine ($\eta_{o}\rightarrow 0.5$), nearly matching the Carnot engine efficiency ($\eta_{c}=0.5$). On the other hand, when the system operates as a refrigerator, both the Carnot and Otto cycles exhibit nearly the same thermal efficiency, as given by (18). However, for higher magnetic fields, the Otto efficiency decreases.
Figure 5: Thermal efficiency of Carnot and Otto engine as a function of magnetic field, considering $U=10$, $V=5.4$, $\mu=15.8$. Particularly here we fixed the $T_{h}=0.56$ and $T_{l}=0.28$.
[…]
Figure 6: Heat absorbed ($Q_{in}$), heat released ($Q_{out}$), and work performed ($W_{net}$) as functions of the magnetic field, assuming $U=10$, $V=5.3$, and $\mu=15.8$. Specifically, here we have fixed $T_{h}=0.43$ and $T_{l}=0.19$.
[…]
Figure 7: Thermal efficiency of Carnot and Otto engine as a function of magnetic field, considering $U=10$, $V=5.3$, $\mu=15.8$. Particularly here we fixed the $T_{h}=0.43$ and $T_{l}=0.19$.
The near-isothermal behavior observed during the isomagnetic process where the Otto cycle closely mimics a Carnot engine underscores the potential for optimized thermal control. Such behavior, characterized by a pronounced magnetocaloric effect, may provide a strategic advantage in designing quantum engines with efficiency approaching the Carnot limit.
Of course, any engine can be manipulated in accordance with the fundamental Carnot theorem of the second law of thermodynamics to achieve Carnot efficiency. However, our main result demonstrates a natural way in which the Otto engine closely mimics the Carnot engine near the anomalous region. This implies that the Otto engine could approach the performance of an almost perfect machine.
V Conclusion
This study investigates the theoretical application of the one-dimensional EHM in the atomic limit, a fundamental physical property initially studied in the 1970s [31, 32, 33]. Subsequent studies, particularly by Glocke [29] and Mancini [35], revealed its anomalous behavior, which has recently been reexamined to uncover finite-temperature pseudo-transitions in certain 1D models [43]. These pseudo-transitions occur in specific regimes, offering intriguing insights into quantum many-body systems.
Given the growing interest in quantum many-body machines and their technological potential, we explored the Carnot and Otto machines with the 1D EHM in the atomic limit as the working substance. Our primary focus was the anomalous region, where pseudo-transitions occur, particularly the transition between the alternating pair (AP) and paramagnetic (PM) phases [43]. Near the pseudo-critical region, our analysis shows that the Otto engine closely mimics the Carnot engine, with its performance approaching that of an ideal machine. While both engines operate through a four-stroke cycle, there are notable distinctions: the Carnot cycle involves two adiabatic and two isothermal processes performed quasi-statically to maintain fixed temperatures, whereas the Otto engine uses two isomagnetic processes (fixed magnetic field) where temperatures are allowed to vary. This adaptability gives the Otto engine an edge in practical scenarios, as it combines high efficiency with operational simplicity.
Although the second law of thermodynamics prohibits any engine from exceeding Carnot efficiency, our findings demonstrate that, within the anomalous region, the Otto machine mimics quite well Carnot machine. Note that this process is not occurring in phase transition region but in an anomalous region only. This underscores the potential of many-body systems in enhancing quantum thermodynamic performance.
It is well known that engines operating between the same hot and cold reservoirs can achieve similar efficiencies, but their practical effectiveness often depends on physical constraints. For instance, Carnot cycles, while theoretically ideal, require long time intervals due to the quasi-static nature of their adiabatic processes. Such long durations can limit their power output. In contrast, the Otto cycle offers a practical advantage by enabling faster cycles, making it more efficient for generating work within shorter timeframes. This highlights the importance of analyzing and comparing different engine cycles, as we have done here, to identify trade-offs between efficiency and power output, similarly that was explored in finite-time thermodynamics[62].
Figs. 1 and 2 illustrate that the pseudo-transition region offers a unique thermodynamic regime with smooth entropy variation, enabling efficient quantum thermodynamic cycles. Unlike true critical points, which require fine-tuning or result in divergences, the pseudo-transition region provides stable thermodynamic control. These findings suggest that engineered quantum materials exhibiting pseudo-transitions could be promising candidates for quantum heat engines and other thermodynamic applications.
While our study focuses on the 1D EHM in the atomic limit, the principles and findings can likely extend to other systems exhibiting similar anomalous properties. This paves the way for exploring alternative theoretical engines and identifying unique advantages offered by different working substances. The anomalous behavior observed in this model is just one application, and future work could explore other characteristics of such systems, potentially revealing new and unexpected behaviors that challenge conventional thermodynamic paradigms.
Although our analysis is confined to a one-dimensional framework, the insights gained offer a conceptual stepping stone for extending these ideas to higher dimensions. In particular, the pronounced response functions observed here could qualitatively guide experimental investigations in quasi-1D or even 2D/3D systems, where similar microscopic reconfigurations might occur under finite-size effects.
Acknowledgements.
O.R., M. R. and S. M. de S. thank Brazilian agencies FAPEMIG and CNPq for partial financial support.
Appendix A Carnot and Otto Process
This appendix provides a comprehensive overview of both Carnot and Otto engines, detailing their principles, operation, and key characteristics.
A.1 Quantum many-body Carnot machine cycle
Here, we explore the typical operation of a reversible Carnot engine, as illustrated in Fig. 8.
Figure 8: A schematic representation of a quantum many-body Carnot machine, assuming $T_{l}=T_{p_{1}}$ and $T_{h}=T_{p_{2}}$.
  1. 1.
    The process $a\rightarrow b$ is adiabatic, meaning no heat is exchanged ($Q_{ab}=0$). According to a fundamental thermodynamic result, the work done during this process is given by:
    $W_{ab}=$ $-\Delta{\cal E}_{ab}=-({\cal E}_{b}-{\cal E}_{a})$
    $=$ ${\cal E}(T{}_{l},h_{0})-{\cal E}(T_{h},h^{\prime}_{1}),$ (21)
    where $h_{0}$ can be found from the adiabatic condition $\mathcal{S}(T_{l},h^{\prime}_{0})=\mathcal{S}(T_{h},h_{1})=\mathcal{S}_{1}$.
  2. 2.
    The process $b\rightarrow c$ is an isothermal process, with $\mathrm{d}T=0$. The heat absorbed during this process is
    $Q_{bc}=$ $T_{h}\left[\mathcal{S}(T_{h},h_{1})-\mathcal{S}(T_{h},h^{\prime}_{1})\right]$
    $=$ $T_{h}\left[\mathcal{S}(T_{h},h_{1})-\mathcal{S}(T_{l},h_{0})\right]$ (22)
    $=$ $T_{h}\left(\mathcal{S}_{1}-\mathcal{S}_{0}\right)=Q_{in}>0,$ (23)
    here $\mathcal{S}(T_{l},h_{0})=\mathcal{S}_{0}$.
  3. 3.
    The process $c\rightarrow d$ is adiabatic, so no heat is exchanged $Q_{cd}=0$. The work done during this process becomes
    $W_{cd}=$ $-\Delta{\cal E}_{cd}=-({\cal E}_{d}-{\cal E}_{c})$
    $=$ ${\cal E}(T{}_{h},h_{1})-{\cal{\cal E}}(T_{l},h^{\prime}_{0}).$ (24)
  4. 4.
    The process $d\rightarrow a$ is isothermal, with ${\rm d}T=0$. The heat released during this process is expressed by
    $Q_{da}=$ $T_{l}\left[\mathcal{S}(T_{l},h_{0})-\mathcal{S}(T_{l},h^{\prime}_{0})\right]$ (25)
    $=$ $T_{l}\left[\mathcal{S}(T_{l},h_{0})-\mathcal{S}(T_{h},h_{1})\right]$ (26)
    $=$ $T_{l}\left(\mathcal{S}_{0}-\mathcal{S}_{1}\right)=Q_{out}<0.$ (27)
    Therefore, the work done during closed reversible cyclic becomes
    $W_{net}=$ $Q_{bc}+Q_{da}.$
    The corresponding heat absorbed when the system’s entropy changes from $\mathcal{S}_{0}$ to $\mathcal{S}_{1}$ is given by
    $Q_{in}=Q_{bc}=T_{h}\left(\mathcal{S}_{1}-\mathcal{S}_{0}\right).$ (28)
When the system operates as an engine, the corresponding heat engine efficiency is described by
$\eta_{c}=$ $\frac{W_{net}}{Q_{bc}}=1+\frac{Q_{da}}{Q_{bc}}=1-\frac{T_{l}}{T_{h}}=1-\frac{T_{p_{1}}}{T_{p_{2}}}.$ (29)
In contrast, when the system operates in other engine modes, the corresponding thermal efficiency is defined in Table 1.
A.2 Quantum many-body Otto machine cycle
Now, let us analyze a typical reversible Otto engine cycle. A schematic representation of the process is depicted in Fig. 9.
Figure 9: Schematic representation of quantum many-body Otto machine, assuming $T_{l}<T^{\prime}_{l}$ with $T_{l}<T_{p_{1}}$ and $T^{\prime}_{h}<T_{h}$ with $T^{\prime}_{h}<T_{p_{2}}$.
  1. 1.
    The process $a\rightarrow b$ is adiabatic, meaning no heat is transferred $Q_{ab}=0$. Let us assume a constant entropy, $\mathcal{S}(T_{l},h_{0})=\mathcal{S}(T^{\prime}_{h},h_{1})=\mathcal{S}_{0}$. Using the first law of thermodynamics, the work done during this process is given by
    $W_{ab}=$ $-\Delta{\cal E}_{ab}=-({\cal E}_{b}-{\cal E}_{a})$
    $=$ $\mathcal{E}(T_{l},h_{0})-\mathcal{E}(T^{\prime}_{h},h_{1}),$ (30)
  2. 2.
    The process $b\rightarrow c$ is an isomagnetic (constant magnetic field) process, thus ${\rm d}h=0$. This implies the work is null. Therefore, the heat transferred becomes
    $Q_{bc}=$ $\Delta\mathcal{E}_{bc}=\mathcal{E}(T_{h},h_{1})-\mathcal{E}(T^{\prime}_{h},h_{1})=Q_{in}>0.$ (31)
  3. 3.
    The process $c\rightarrow d$ follows an adiabatic process, which means there is no heat transferred $Q_{cd}=0$. By fixing constant entropy at $\mathcal{S}(T_{h},h_{1})=\mathcal{S}(T^{\prime}_{l},h_{0})=\mathcal{S}_{1}$, only work is performed during this process
    $W_{cd}=$ $-\Delta{\cal E}_{cd}=-({\cal E}_{d}-{\cal E}_{c})$
    $=$ $\mathcal{E}(T_{h},h_{1})-\mathcal{E}(T^{\prime}_{l},h_{0}).$ (32)
  4. 4.
    The process $d\rightarrow a$ is an isomagnetic process, meaning ${\rm d}h=0$. This implies that no work is performed during this process. Therefore, when energy changes occur, only heat is transferred
    $Q_{da}=$ $\Delta\mathcal{E}_{da}=\mathcal{E}(T_{l},h_{0})-\mathcal{E}(T^{\prime}_{l},h_{0})=Q_{out}<0.$ (33)
Appendix B Otto Machine efficiency
In this Appendix, we analyze the efficiency of the Otto machine around the pseudo-transition region. At the pseudo-critical temperature, the specific heat exhibits a very sharp and narrow peak. Therefore, the Taylor expansion around the pseudo-critical temperature $T_{p}$ can be written as follows
$C=C_{p}+\frac{1}{2}C_{p}^{(2)}(T-T_{p})^{2}+\cdots$ (34)
Here, we have that $C_{p}^{(1)}=\left(\frac{\partial C}{\partial T}\right)_{T_{p}}=0$, which corresponds to the maximum of the specific heat (not to be confused with the specific heat at constant pressure), while $C_{p}^{(2)}=\left(\frac{\partial^{2}C}{\partial T^{2}}\right)_{T_{p}}<0$. On the other hand, the following relation holds
$\left(\frac{\partial\mathcal{S}}{\partial T}\right)=\frac{C}{T}.$ (35)
Integrating around $T\rightarrow T_{p}$, we can obtain the entropy as follows
$\intop_{\mathcal{S}_{p}}^{\mathcal{S}}{\rm d}\mathcal{S}=$ $\intop_{T_{p}}^{T}\frac{1}{T}\left[C_{p}+\frac{1}{2}C_{p}^{(2)}(T-T_{p})^{2}\right]{\rm d}T$
$\mathcal{S}-\mathcal{S}_{p}=$ $C_{p}\ln\left(\frac{T}{T_{p}}\right)+\frac{1}{4}C_{p}^{(2)}\left(T-T_{p}\right)\left(T-3T_{p}\right)$
$+\frac{1}{2}C_{p}^{(2)}\ln\left(\frac{T}{T_{p}}\right),$ (36)
$\mathcal{S}-\mathcal{S}_{p}=$ $\left(C_{p}+\frac{T_{p}^{2}}{2}C_{p}^{(2)}\right)\ln\left(\frac{T}{T_{p}}\right)$
$+\frac{1}{4}C_{p}^{(2)}\left(T-T_{p}\right)\left(T-3T_{p}\right).$ (37)
If we wish to express the temperature analytically from equation (36), we encounter a transcendental equation that must be solved numerically. However, for our purposes, we will consider only the first-order approximation as the relevant term, from which the temperature can be expressed analytically. Therefore, from (36), we obtain the following simple expression
$\mathcal{S}=\mathcal{S}_{p}+C_{p}\ln\left(\frac{T}{T_{p}}\right).$ (38)
Now, we can express the temperature as a function of entropy
$T=$ $T_{p}{\rm e}^{\left(\mathcal{S}-\mathcal{S}_{p}\right)/C_{p}}.$
From these relations, we can express the cold temperatures as follows
$T_{l}=$ $T_{p_{1}}{\rm e}^{\left(\mathcal{S}_{0}-\mathcal{S}_{p_{1}}\right)/C_{p_{1}}}$
$T^{\prime}_{l}=$ $T_{p_{1}}{\rm e}^{\left(\mathcal{S}_{1}-\mathcal{S}_{p_{1}}\right)/C_{p_{1}}.}$
Substituting these into the heat released $Q_{da}$ during the d-a process of the Otto cycle (as given in equation (33)), we obtain:
$Q_{da}=$ $C_{p_{1}}T_{p_{1}}\left[{\rm e}^{\mathcal{S}_{0}/C_{p_{1}}}-{\rm e}^{\mathcal{S}_{1}/C_{p_{1}}}\right]{\rm e}^{-\mathcal{S}_{p_{1}}/C_{p_{1}}}$
$\approx$ $T_{p_{1}}\left(\mathcal{S}_{0}-\mathcal{S}_{1}\right)\left[1+\tfrac{1}{2C_{p_{1}}}\left(\mathcal{S}_{0}+\mathcal{S}_{1}-2\mathcal{S}_{p_{1}}\right)\right].$ (39)
Similarly, we can express the entropies for two fixed hot temperature reservoirs as follows
$T^{\prime}_{h}=$ $T_{p_{2}}{\rm e}^{\left(\mathcal{S}_{0}-\mathcal{S}_{p_{2}}\right)/C_{p_{2}}}$
$T_{h}=$ $T_{p_{2}}{\rm e}^{\left(\mathcal{S}_{1}-\mathcal{S}_{p_{2}}\right)/C_{p_{2}}}.$
Analogous to the previous result, the heat absorbed $Q_{bc}$ during the b-c process (as described by equation (31)) becomes
$Q_{bc}=$ $C_{p_{2}}T_{p_{2}}\left[{\rm e}^{\mathcal{S}_{1}/C_{p_{2}}}-{\rm e}^{\mathcal{S}_{0}/C_{p_{2}}}\right]{\rm e}^{-\mathcal{S}_{p_{2}}/C_{p_{2}}}$
$\approx$ $T_{p_{2}}\left(\mathcal{S}_{1}-\mathcal{S}_{0}\right)\left[1+\tfrac{1}{2C_{p_{2}}}\left(\mathcal{S}_{0}+\mathcal{S}_{1}-2\mathcal{S}_{p_{2}}\right)\right].$ (40)
Therefore, the ratio between heat released and absorbed is
$\frac{Q_{da}}{Q_{bc}}\approx$ $-\frac{T_{p_{1}}}{T_{p_{2}}}\frac{\left(1+\tfrac{1}{2C_{p_{1}}}\left(\mathcal{S}_{0}+\mathcal{S}_{1}-2\mathcal{S}_{p_{1}}\right)\right)}{\left(1+\tfrac{1}{2C_{p_{2}}}\left(\mathcal{S}_{0}+\mathcal{S}_{1}-2\mathcal{S}_{p_{2}}\right)\right)},$
$\approx$ $-\frac{T_{p_{1}}}{T_{p_{2}}}\left(1-\tfrac{1}{2C_{p_{2}}}\left(\mathcal{S}_{0}+\mathcal{S}_{1}-2\mathcal{S}_{p_{2}}\right)\right),$
$\approx$ $-\frac{T_{p_{1}}}{T_{p_{2}}}+\frac{T_{p_{1}}}{T_{p_{2}}C_{p_{2}}}\left(\frac{\mathcal{S}_{0}+\mathcal{S}_{1}}{2}-\mathcal{S}_{p_{2}}\right),$ (41)
since $C_{p_{1}}$ is considerable larger than $C_{p_{2}}$, the last term is positive. From this result, one can derive the results given in eq.(20).